OpenStudy (anonymous):

Two regular octagons have areas of 48 mm² and x² + 8x mm². What values of x make the octagons congruent? 6 and 8 6 and -8 4 and 12 4 and -12

Directrix (directrix):

If two polygons are congruent, then they have equal areas.

Directrix (directrix):

48 = x^2 + 8x

Directrix (directrix):

x^2 + 8x -48 = 0

Directrix (directrix):

(x +12)(x-4) = 0 Zero Product Property x + 12=0 or x-4 =0 x = ? or x = ?

OpenStudy (anonymous):

x= -12 and x= 4?

Directrix (directrix):

Yes. x is not a side of the polygon so there is no problem with the negative 12. When substituted in for x in x^2 + 8x, it cranks out a positve result.

OpenStudy (anonymous):

awesome thanks for all your help i can now understand and do them for myself now :)